1:-(50) Find: RG ww 15V I vos , Im and the region of goeration Ro km = 60 mA/v? V2V Real IMR 12 / RD 2002 = and the region
Answer #1

2001 IMA + VOD = 1SV. Vos + VGs. kn = Goma/ur , vt from Circuit vas = Vos. Nos > VGs-Vt •Condition for saturation region NuosVas = Vos = 15-20010. Vas 15 - 200x 58.04x173 Vas = 3:392 V. LA trans conductance gmi VGs 58.04X10 3.392 ms = 17.111m (oroms